OpenStudy (anonymous):

Find the volume of the solid generated by revolving the region in the first quadrant bounded by y=x^2+1 and y=0 about the x axis.

OpenStudy (anonymous):

I keep getting 28pi/15 for the answer but apparently it is 8pi/15

OpenStudy (anonymous):

I used the disc method \[\int\limits_{0}^{1}\pi(x^2+1)^2\]

OpenStudy (anonymous):

\[\pi \int\limits_{0}^{1}x^4+2x^2+1\]

OpenStudy (anonymous):

\[\pi(\frac{ x^5 }{ 5 }+\frac{2 x^3 }{ 3 }+x)\]

OpenStudy (anonymous):

\[\pi \left[ \left( \frac{ 1 }{ 5 }+\frac{ 2 }{ 3 }+1 \right)-0 \right]\]

OpenStudy (anonymous):

which was 28pi/15

OpenStudy (anonymous):

Can anyone tell me where I'm going wrong?

OpenStudy (anonymous):

who told you the bounds are from 0 to 1? there aren't in the question you posted

OpenStudy (anonymous):

Am I using the wrong bound? That's the whole question.

OpenStudy (anonymous):

is that the entire question though? because it doesn't state what the x bounds are, just the y bounds

OpenStudy (anonymous):

That is the entire question

OpenStudy (anonymous):

it can't be :-X lol

OpenStudy (anonymous):

x^2 + 1 goes on forever

OpenStudy (anonymous):

It is, haha. I'm studying past final exams and it's a multiple choice question

OpenStudy (anonymous):

makes no sense at all. but you are doing it right. the problem must be the bounds

OpenStudy (anonymous):

hmmm, this is frustrating. I did notice that for the final answer (3/15 + 10/15 + 15/15)=28/15....(3/15+15/15-10/15)=8/15....which is apparently the correct answer

OpenStudy (anonymous):

I think I just may email my professor. It looks like there may be a typo. Maybe it's supposed to be -x^2+1? Then x could be + or -1

OpenStudy (anonymous):

Thank you for looking at it for me

OpenStudy (anonymous):

yes, 1 - x^2 makes a lot of sense. the bounds would be implied for this one and the answer is actually 8pi/15. great find

OpenStudy (anonymous):

I bet that's the problem then. It looks like it got messed up in the printer or something because the equation is all on top of eachother so it's difficult to read.

OpenStudy (anonymous):

ya :)